# From Rows to Reasoning: A Retrieval-Augmented Multimodal Framework for Spreadsheet Understandingthanks: FRTR-Bench dataset: https://github.com/AnmolGulati6/FRTR-bench

### v1

@@ Abstract
Large Language Models (LLMs) struggle to reason over large-scale enterprise spreadsheets containing thousands of numeric rows, multiple linked sheets, and embedded visual content such as charts and receipts. Prior state-of-the-art spreadsheet reasoning approaches typically rely on single-sheet compression or full-context encoding, which limits scalability and fails to reflect how real users interact with complex, multimodal workbooks. We introduce FRTR-Bench, the first large-scale benchmark for multimodal spreadsheet reasoning, comprising 30 enterprise-grade Excel workbooks spanning nearly four million cells and more than 50 embedded images. To address these challenges, we present From Rows to Reasoning (FRTR), an advanced, multimodal retrieval-augmented generation framework that decomposes Excel workbooks into granular row, column, and block embeddings, employs hybrid lexical-dense retrieval with Reciprocal Rank Fusion (RRF), and integrates multimodal embeddings to reason over both numerical and visual information. We tested FRTR on six LLMs, achieving 74% answer accuracy on FRTR-Bench with Claude Sonnet 4.5, a substantial improvement over prior state-of-the-art approaches that reached only 24%. On the SpreadsheetLLM benchmark, FRTR achieved 87% accuracy with GPT-5 while reducing token usage by roughly 50% compared to context-compression methods.
1 Introduction
@@
Spreadsheets remain the foundational medium for analytical work at an enterprise level, and modern workbooks routinely contain thousands of cells, span multiple sheets, and embed diverse media such as charts, receipts, and scanned tables [17, 14]. While Large Language Models (LLMs) promise intuitive natural-language interaction with such documents, directly serializing entire sheets into a prompt is inefficient and brittle. Spreadsheet semantics hinge on spatial layout, structural dependencies, and cross-sheet relationships rather than text alone. Moreover, recent work shows that long-context LLMs suffer positional degradation, as accuracy declines when relevant information appears far from the prompt boundaries [25]. This motivates retrieval-focused approaches that surface small, relevant subsets rather than streaming full spreadsheets into the context window.
@@
A prominent direction by Haoyu Dong et al., SpreadsheetLLM [10], introduced encoding schemes SheetEncoder and SheetCompressor that compress cell content, positional addresses, and formatting to fit within an LLM’s context window. While effective for single-sheet reasoning, such compression assumes all relevant data can fit within a single prompt and fails to scale to multi-sheet or multimodal settings. In practice, enterprise workbooks exceed these boundaries, incorporating thousands of rows, inter-sheet references, and embedded images. Retrieval-augmented generation approaches provide a complementary path: combining database-style retrieval with hybrid lexical–dense retrieval to fetch only the minimal, most relevant evidence for reasoning [6, 18, 19]. Empirical studies demonstrate that retrieval-augmented generation (RAG) can match or surpass long-context models at significantly lower computational cost and latency [22, 13, 3, 18]. Simple rank fusion methods, such as Reciprocal Rank Fusion (RRF), further enhance stability and performance across retrieval modalities [7].
Figure 1: Excel workbooks are decomposed into rows, columns, sliding context windows, and embedded images during indexing. These components are stored in a multimodal hybrid vector database supporting lexical and dense fields. Hybrid lexical–dense retrieval (RRF) fetches relevant evidence for LLM reasoning.
@@
2 Related Work
2.1 LLMs for Tables and Spreadsheets
@@
Early neural models such as TAPAS [15] learned to answer questions directly over tables by encoding cell values jointly with queries, while TUTA [32] extended this paradigm by modeling hierarchical and spatial cues. OmniTab [8] further improved few-shot table QA by combining natural and synthetic supervision. These approaches focus on in-context reasoning within a single table and do not scale to multi-sheet or multimodal workbooks.
@@
More recent work such as SpreadsheetLLM [10] compresses spreadsheet structure and content to fit within LLM context limits, achieving notable gains on single-sheet reasoning. Complementary methods, including Table-GPT [5] and TableLLM [4], improve general table manipulation and code generation capabilities. However, these approaches are not capable of handling cross-sheet and multimodal reasoning. In contrast, FRTR emphasizes hybrid lexical-dense retrieval and multimodal embedding [11, 34, 1, 29, 28].
2.2 Long Context, Retrieval, and Hybrid Search
@@
Despite advances in long-context modeling, evidence retrieval remains a bottleneck: simply extending context windows does not guarantee that models attend effectively to mid-range information [25]. Retrieval-augmented generation approaches address this by selecting relevant evidence before generation, improving both performance and interpretability [21, 13]. Hybrid lexical-dense retrieval, fused through methods such as Reciprocal Rank Fusion (RRF) [6, 7], offers robustness and simplicity across tasks. However, basic chunking and retrieval methodologies do not take into account the complicated spatial relationships in spreadsheets, where logic can apply across rows, columns, and even sheets.
2.3 Multimodal Spreadsheet Understanding and Document VQA
@@
Spreadsheets often convey semantics through visual layout, including bold headers, merged cells, and shaded totals, as well as through embedded images such as charts or scanned receipts. Vision–language models (VLMs) still struggle to interpret such spatial and structural cues [24, 30, 16, 23]. Parallel research in chart and document VQA (e.g., ChartQA [26], DocVQA [27]) highlights the challenges of numerically grounded visual reasoning. These methods typically treat visual and textual modalities in isolation, without integrating multimodal embeddings that unify text, layout, and image semantics.
2.4 Benchmarks for Spreadsheet Interaction
@@
Beyond table QA, recent benchmarks target spreadsheet interaction directly. SpreadsheetLLM [10] performs reasoning on encoded spreadsheets, while SpreadsheetBench [20] introduces realistic, forum-based spreadsheet questions. InstructExcel [33] focuses on natural-language-to-OfficeScript translation. No existing benchmarks offer assessment involving cross-sheet reasoning, multimodal queries, and large-scale evaluation, reflecting the real-world challenges of enterprise workbooks.
3 Framework Overview
3.1 Motivation and Design Goals
@@
Real-world Excel workbooks exhibit three intertwined challenges: scale, multimodality, and cross-sheet dependency. Workbooks can span thousands of rows across multiple worksheets, include embedded charts or receipts, and encode meaning through layout and formatting. Existing approaches either compress entire files into an LLM’s context, trading precision for token cost, or rely on text-only embeddings that ignore visual cues.
@@
  1. 1.
    Scalability: Decompose workbooks into granular, retrievable units such as rows, columns, blocks, and images to bypass context limits.
  2. 2.
@@
3.2 Architecture Overview
As shown in Figure 1, FRTR operates in three stages: Retrieve, Verify, and Compose. The pipeline is formalized in Algorithm 1. During ingestion, Excel workbooks are decomposed into rows, columns, sliding windows, and embedded images. Each unit is serialized and embedded using a multimodal encoder (e.g., Amazon Titan Multimodal) and indexed in a hybrid vector database [2]. At query time, FRTR retrieves the most relevant units using hybrid lexical–dense search and provides the fused results as evidence for LLM reasoning. This architecture reduces context size while preserving interpretability and scalability.
3.3 Retrieval and Fusion Layer
Each spreadsheet unit is indexed as a structured document with lexical, vector, and metadata fields. FRTR performs two parallel retrievals: BM25 for lexical matching and cosine similarity for dense embeddings, and fuses results via Reciprocal Rank Fusion (RRF) [6, 31].
@@ 3.3 Retrieval and Fusion Layer
where $k=60$ stabilizes low-ranked items. FRTR uses $K_{v}=K_{s}=20$, returning the top-10 fused items annotated with provenance. This rank-based fusion avoids score normalization issues, ensuring robust, interpretable retrieval across modalities.
@@
3.4 Multimodal Embedding Layer
@@
These are embedded via the text encoder. Image units (charts, receipts, diagrams) are embedded using the vision branch of the same model, ensuring a shared latent space. This allows queries such as “Q4 revenue trends” to retrieve both numerical data and corresponding charts. Generated embeddings $\mathbf{v}_{u}$ are stored in a unified vector field that holds both text and image representations. This unified field enables cross-modal retrieval within a single index, maintaining simplicity and generality.
3.5 Reasoning and Output Generation
@@
This format ensures both interpretability and machine-readability, facilitating downstream automation and audit compliance. The complete prompt template used in this study is provided in Section Appendix C. FRTR itself does not execute formulas or modify spreadsheets, maintaining human oversight while ensuring verifiable, grounded reasoning.
4 FRTR-Bench: Benchmark for Spreadsheet Reasoning
4.1 Dataset Construction
Existing benchmarks like SpreadsheetLLM [10] and SpreadsheetBench [20] focus on single-sheet, text-only reasoning. To evaluate cross-sheet and multimodal reasoning at scale, we introduce FRTR-Bench, a dataset of 30 multimodal Excel workbooks across diverse domains. Similar to how Jackal introduced an execution-based, real-world benchmark for text-to-JQL tasks to evaluate LLMs under production-like conditions, we design FRTR-Bench to capture enterprise spreadsheet reasoning at realistic scale and modality [12]. Each workbook stresses three dimensions: (1) Scale: up to 210,000 rows per workbook, far beyond typical LLM context; (2) Multimodality: embedded PNG images such as charts, receipts, and dashboards; and (3) Cross-sheet complexity: multi-sheet references (e.g., =SUM(Sheet1!B2:B100, Sheet2!C5:C50)).
@@ 4.2 Evaluation Protocol
4.3 Benchmark Statistics and Examples
Table 1: FRTR-Bench dataset statistics summarizing workbook, sheet, and multimodal content scale.
@@ 4.3 Benchmark Statistics and Examples
For example, in frtr_0003_consolidation.xlsx (Hard, 293K cells):
Question: What is the consolidated operating income for Region EMEA in FY2024? ReasoningType: cross-sheet Provenance: Consolidation!E47,  =SUM(EMEA_Sales!D:D, EMEA_Ops!E:E) Answer: \$12,450,000
A full-context LLM would need to serialize hundreds of thousands of cells, far beyond feasible context limits, while FRTR retrieves fewer than 10 relevant chunks (<5K tokens) to reason effectively. FRTR-Bench thus provides a rigorous, multimodal, and scalable benchmark for retrieval-augmented spreadsheet reasoning, capturing the challenges of real-world analytical workflows.
5 Evaluation
@@
All experiments use the same evaluation criteria defined in Section 4: Answer Accuracy, Latency (s), and Mean Tokens.
@@
All experiments evaluate six LLMs spanning both small and large, and open- and closed-weight architectures: GPT-4o, GPT-5, Claude 4.5 Sonnet, Gemini 2.5 Pro, LLaMA Maverick-17B Instruct, and GPT-OSS-120b.
Table 2: Performance comparison of FRTR and SpreadsheetLLM across multiple LLMs on FRTR-Bench. The SpreadsheetLLM algorithm was modified to parallelize its compression pipeline and include all sheets as context[9].
Technique Model Answer Accuracy Mean Tokens Latency (s)
FRTR (ours) GPT-4o 0.49 7,690.91 5.04
GPT-5 0.73 7,690.91 15.50
Sonnet-4.5 0.74 7,690.91 11.71
Gemini-2.5-Pro 0.67 7,690.91 26.90
LLaMA-Maverick-17B-instruct 0.56 7,690.91 2.42
GPT-OSS-120b 0.51 7,690.91 5.75
SpreadsheetLLM GPT-4o 0.06 12,744.64 1.14
GPT-5 0.18 12,744.64 18.10
Sonnet-4.5 0.13 12,744.64 8.80
Gemini-2.5-Pro 0.24 12,744.64 33.35
LLaMA-Maverick-17B-instruct 0.23 12,744.64 1.15
GPT-OSS-120b 0.21 12,744.64 8.80
@@
Figure 2: Performance comparison by difficulty level across methods on FRTR-Bench.
5.1 Experiment 1 - FRTR-Bench Evaluation
5.1.1 Goal

### v2

@@ Abstract.
Large Language Models (LLMs) struggle to reason over large-scale enterprise spreadsheets containing thousands of numeric rows, multiple linked sheets, and embedded visual content such as charts and receipts. Prior state-of-the-art spreadsheet reasoning approaches typically rely on single-sheet compression or full-context encoding, which limits scalability and fails to reflect how real users interact with complex, multimodal workbooks. We introduce FRTR-Bench, the first large-scale benchmark for multimodal spreadsheet reasoning, comprising 30 enterprise-grade Excel workbooks spanning nearly four million cells and more than 50 embedded images. To address these challenges, we present From Rows to Reasoning (FRTR), an advanced, multimodal retrieval-augmented generation framework that decomposes Excel workbooks into granular row, column, and block embeddings, employs hybrid lexical-dense retrieval with Reciprocal Rank Fusion (RRF), and integrates multimodal embeddings to reason over both numerical and visual information. We tested FRTR on six LLMs, achieving 74% answer accuracy on FRTR-Bench with Claude Sonnet 4.5, a substantial improvement over prior state-of-the-art approaches that reached only 24%. On the SpreadsheetLLM benchmark, FRTR achieved 87% accuracy with GPT-5 while reducing token usage by roughly 50% compared to direct serialization methods.
Large Language Models, Retrieval-Augmented Generation (RAG), Multimodal Data Retrieval, Scalable Numerical Reasoning
conference: ;
1. Introduction
@@
Spreadsheets remain the foundational medium for analytical work at an enterprise level, and modern workbooks routinely contain thousands of cells, span multiple sheets, and embed diverse media such as charts, receipts, and scanned tables (Kandel et al., 2012; Hermans and Murphy-Hill, 2015). While Large Language Models (LLMs) promise intuitive natural-language interaction with such documents, directly serializing entire sheets into a prompt is inefficient and brittle. Spreadsheet semantics hinge on spatial layout, structural dependencies, and cross-sheet relationships rather than text alone. Moreover, recent work shows that long-context LLMs suffer positional degradation, as accuracy declines when relevant information appears far from the prompt boundaries (Liu et al., 2023). This motivates retrieval-focused approaches that surface small, relevant subsets rather than streaming full spreadsheets into the context window.
Refer to caption Scalability comparison showing prompt token usage vs spreadsheet rows. Without retrieval, tokens grow linearly and exceed context limits. With FRTR, tokens remain constant at approximately 8K regardless of data size.
Figure 1. Token scaling comparison: naive full-context serialization exceeds context limits as data grows, while FRTR maintains constant token usage ($\sim$8K) via selective retrieval, enabling scalable enterprise reasoning.
@@
A prominent direction is SpreadsheetLLM (Dong et al., 2025b), which introduces encoding schemes SheetEncoder and SheetCompressor that compress cell content, positional addresses, and formatting to fit within an LLM’s context window. While effective for single-sheet reasoning, such compression assumes all relevant data can fit within a single prompt and fails to scale to multi-sheet or multimodal settings. In practice, enterprise workbooks exceed these boundaries, incorporating thousands of rows, inter-sheet references, and embedded images. Retrieval-augmented generation approaches provide a complementary path: combining database-style retrieval with hybrid lexical–dense retrieval to fetch only the minimal, most relevant evidence for reasoning (Cormack et al., 2009; Karpukhin et al., 2020; Khattab and Zaharia, 2020). Empirical studies demonstrate that retrieval-augmented generation (RAG) can match or surpass long-context models at significantly lower computational cost and latency (Lewis et al., 2021b; Gao et al., 2024; Borgeaud et al., 2022; Karpukhin et al., 2020). Simple rank fusion methods, such as Reciprocal Rank Fusion (RRF), further enhance stability and performance across retrieval modalities (Corporation, 2025).
@@
2. Related Work
2.1. LLMs for Tables and Spreadsheets
@@
Early neural models such as TAPAS (Herzig et al., 2020) learned to answer questions directly over tables by encoding cell values jointly with queries, while TUTA (Wang et al., 2021) extended this paradigm by modeling hierarchical and spatial cues. OmniTab (Jiang et al., 2022) further improved few-shot table QA by combining natural and synthetic supervision. These approaches focus on in-context reasoning within a single table and do not scale to multi-sheet or multimodal workbooks.
@@
More recent work such as SpreadsheetLLM (Dong et al., 2025b) compresses spreadsheet structure and content to fit within LLM context limits, achieving notable gains on single-sheet reasoning. Complementary methods, including Table-GPT (Zha et al., 2023) and TableLLM (Chen, 2023), improve general table manipulation and code generation capabilities. However, these approaches are not capable of handling cross-sheet and multimodal reasoning. In contrast, FRTR emphasizes hybrid lexical-dense retrieval and multimodal embedding (Drushchak et al., 2025; Zhang et al., 2025; Abootorabi et al., 2025; Rackauckas, 2024; Mei et al., 2025).
Refer to caption Architecture diagram showing the FRTR framework: Excel workbooks are decomposed into rows, columns, sliding context windows, and embedded images. These are indexed into a multimodal hybrid vector database. At query time, hybrid retrieval using RRF fetches relevant chunks which are passed to an LLM for reasoning.
Figure 2. Excel workbooks are decomposed into rows, columns, sliding context windows, and embedded images during indexing. These components are stored in a multimodal hybrid vector database supporting lexical and dense fields. Hybrid lexical–dense retrieval (RRF) fetches relevant evidence for LLM reasoning.
2.2. Long Context, Retrieval, and Hybrid Search
@@
Despite advances in long-context modeling, evidence retrieval remains a bottleneck: simply extending context windows does not guarantee that models attend effectively to mid-range information (Liu et al., 2023). Retrieval-augmented generation approaches address this by selecting relevant evidence before generation, improving both performance and interpretability (Lewis et al., 2021a; Gao et al., 2024). Hybrid lexical-dense retrieval, fused through methods such as Reciprocal Rank Fusion (RRF) (Cormack et al., 2009; Corporation, 2025), offers robustness and simplicity across tasks. However, basic chunking and retrieval methodologies do not take into account the complicated spatial relationships in spreadsheets, where logic can apply across rows, columns, and even sheets.
2.3. Multimodal Spreadsheet Understanding and Document VQA
@@
Spreadsheets often convey semantics through visual layout, including bold headers, merged cells, and shaded totals, as well as through embedded images such as charts or scanned receipts. Vision–language models (VLMs) still struggle to interpret such spatial and structural cues (Radford et al., 2021; Jia et al., 2021; Li et al., 2023). Parallel research in chart and document VQA (e.g., ChartQA (Masry et al., 2022), DocVQA (Mathew et al., 2021)) highlights the challenges of numerically grounded visual reasoning. These methods typically treat visual and textual modalities in isolation, without integrating multimodal embeddings that unify text, layout, and image semantics.
2.4. Benchmarks for Spreadsheet Interaction
@@
Beyond table QA, recent benchmarks target spreadsheet interaction directly. SpreadsheetLLM (Dong et al., 2025b) performs reasoning on encoded spreadsheets, while SpreadsheetBench (Ma et al., 2024) introduces realistic, forum-based spreadsheet questions. InstructExcel (Payan et al., 2023) focuses on natural-language-to-OfficeScript translation. No existing benchmarks assess cross-sheet reasoning, multimodal queries, and large-scale evaluation together, reflecting the real-world challenges of enterprise workbooks.
3. Framework Overview
3.1. Motivation and Design Goals
@@
Real-world Excel workbooks exhibit three intertwined challenges: scale, multimodality, and cross-sheet dependency. Workbooks can span thousands of rows across multiple worksheets, include embedded charts or receipts, and encode meaning through layout and formatting. Existing approaches either compress entire files into an LLM’s context, trading precision for token cost, or rely on text-only embeddings that ignore visual cues. Although some enterprise spreadsheets contain irregular or weakly aligned layouts, FRTR does not assume well-formed tables; its use of overlapping row-, column-, and window-level units allows relevant evidence to be retrieved even when logical groupings are not strictly aligned with grid boundaries.
@@
  1. (1)
    Scalability: Decompose workbooks into granular, retrievable units such as rows, columns, blocks, and images to bypass context limits, deduplicating merged cells and representing both formulas and calculated values.
  2. (2)
@@
3.2. Architecture Overview
As shown in Figure 2, FRTR operates in three stages: Retrieve, Verify, and Compose. The pipeline is formalized in Algorithm 1. During ingestion, Excel workbooks are decomposed into rows, columns, sliding windows, and embedded images (we investigate the impact of different chunk types in Section 5.3). Each unit is serialized and embedded using a multimodal encoder (e.g., Amazon Titan Multimodal) and indexed in a hybrid vector database (Amazon Web Services, Inc., 2025). At query time, FRTR retrieves the most relevant units using hybrid lexical–dense search and provides the fused results as evidence for LLM reasoning. This architecture reduces context size while preserving interpretability and scalability.
3.3. Retrieval and Fusion Layer
Each spreadsheet unit is indexed as a structured document with lexical, vector, and metadata fields. FRTR performs two parallel retrievals: BM25 for lexical matching and cosine similarity for dense embeddings, and fuses results via Reciprocal Rank Fusion (RRF) (Cormack et al., 2009; Robertson and Zaragoza, 2009).
@@ 3.3. Retrieval and Fusion Layer
where $k=60$ stabilizes low-ranked items. FRTR uses $K_{v}=K_{s}=20$, returning the top-10 fused items annotated with provenance. This rank-based fusion avoids score normalization issues, ensuring robust, interpretable retrieval across modalities. RRF is used because it provides the most robust retrieval across both paraphrased queries and exact identifier lookups common in spreadsheet QA (see 5.3).
@@
3.4. Multimodal Embedding Layer
@@
These are embedded via the text encoder. Image units such as charts, receipts, and diagrams are embedded using the vision branch of the same model, ensuring a shared latent space. In our implementation, we encode images as base64 and pass them to the Titan Image-v1 embeddings API to generate embeddings. This allows queries such as ”Q4 revenue trends” to retrieve both numerical data and corresponding charts. Generated embeddings $\mathbf{v}_{u}$ are stored in a unified vector field that holds both text and image representations. This unified field enables cross-modal retrieval within a single index, maintaining simplicity and generality.
3.5. Reasoning and Output Generation
@@
This format ensures both interpretability and machine-readability, facilitating downstream automation and audit compliance. The complete prompt template used in this study is provided in Appendix C. FRTR itself does not execute formulas or modify spreadsheets, maintaining human oversight while ensuring verifiable, grounded reasoning.
4. FRTR-Bench: Benchmark for Spreadsheet Reasoning
4.1. Dataset Construction
Existing benchmarks like SpreadsheetLLM (Dong et al., 2025b) and SpreadsheetBench (Ma et al., 2024) focus on single-sheet, text-only reasoning. To evaluate cross-sheet and multimodal reasoning at scale, we introduce FRTR-Bench, a dataset of 30 multimodal Excel workbooks across diverse domains. Similar to how Jackal introduced an execution-based, real-world benchmark for text-to-JQL tasks to evaluate LLMs under production-like conditions, we design FRTR-Bench to capture enterprise spreadsheet reasoning at realistic scale and modality (Frank et al., 2025). Each workbook stresses three dimensions: (1) Scale: up to 210,000 rows per workbook, far beyond typical LLM context; (2) Multimodality: embedded PNG images such as charts, receipts, and dashboards; and (3) Cross-sheet complexity: multi-sheet references (e.g., =SUM(Sheet1!B2:B100, Sheet2!C5:C50)).
@@ 4.2. Evaluation Protocol
4.3. Benchmark Statistics and Examples
Table 1. FRTR-Bench dataset statistics summarizing workbook, sheet, and multimodal content scale.
@@ 4.3. Benchmark Statistics and Examples
For example, in frtr_0003_consolidation.xlsx (Hard, 49,867 rows):
Question: What is the consolidated operating income for Region EMEA in FY2024?
ReasoningType: cross-sheet
Provenance: Consolidation!E47,
  =SUM(EMEA_Sales!D:D,EMEA_Ops!E:E)
Answer: \$12,450,000
Table 2. Performance comparison on FRTR-Bench. FRTR consistently outperforms SpreadsheetLLM while using 40% fewer tokens. Mean tokens: FRTR 7,691; SpreadsheetLLM 12,745. We implemented the SpreadsheetLLM algorithm as written (Dong et al., 2025a).
FRTR (Ours) SpreadsheetLLM
Model Accuracy Latency (s) Accuracy Latency (s) $\Delta$ Accuracy
GPT-4o 0.49 5.04 0.06 1.14 +0.43
GPT-5 0.73 15.50 0.18 18.10 +0.55
Claude Sonnet 4.5 0.74 11.71 0.13 8.80 +0.61
Gemini 2.5 Pro 0.67 26.90 0.24 33.35 +0.43
LLaMA Maverick-17B 0.56 2.42 0.23 1.15 +0.33
GPT-OSS-120B 0.51 5.75 0.21 8.80 +0.30
A full-context LLM would need to serialize hundreds of thousands of cells, far beyond feasible context limits, while FRTR retrieves fewer than 10 relevant chunks (<<5K tokens) to reason effectively. FRTR-Bench thus provides a rigorous, multimodal, and scalable benchmark for retrieval-augmented spreadsheet reasoning, capturing the challenges of real-world analytical workflows.
5. Evaluation
@@
All experiments use the same evaluation criteria defined in Section 4: Answer Accuracy, Latency (s), and Mean Tokens. Latency refers to the inference time of the LLM.
@@
All experiments evaluate six LLMs spanning both small and large, and open- and closed-weight architectures: GPT-4o, GPT-5, Claude Sonnet 4.5, Gemini 2.5 Pro, LLaMA Maverick-17B Instruct, and GPT-OSS-120b. All LLMs and embedding models were accessed via APIs using standard pricing. While FRTR incurs additional embedding calls during indexing, embedding models are substantially cheaper than LLM inference, so the overall cost is dominated by prompt tokens passed to the LLM and remains lower than full-context baselines.
Baselines.
We compare FRTR against two baseline approaches:
  1. (1)
  2. (2)
Refer to caption Line chart comparing peak accuracy of FRTR and SpreadsheetLLM methods across spreadsheets with varying total row counts. FRTR consistently outperforms SpreadsheetLLM, with the gap widening as spreadsheet size increases.
@@
Figure 3. Peak performance comparison by total rows across methods on FRTR-Bench.
5.1. Experiment 1 - FRTR-Bench Evaluation
5.1.1. Goal